OpenStudy (anonymous):

A rectangular plot of land is 100 feet long and 50 feet wide. How long is the walkway along the diagonal? Round to the nearest foot. A. 75 feet B. 87 feet C. 112 feet D. 150 feet

OpenStudy (anonymous):

You're bisecting the rectangle diagonally. Therefore, you have a typical right triangle problem [(a^2)+(b^2)=(c^2)] You know a and b, and they are interchangeable. So, use (100^2)+(50^2)=(c^2) So, 10000+2500=(c^2) So, 12500=(c^2) So, c=sqrt(12500) So, c=~112 feet

OpenStudy (jdoe0001):